if x= acos pi t find the distance travelled in t= 2.5 seconds?
 

Grade:12th pass

2 Answers

Sumit Majumdar IIT Delhi
8 years ago
The average distance covered would be equal to:
<x_{avg}>=\frac{\int_{0}^{2.5}a\cos{(\pi t)}dt}{2.5}=\frac{[a\sin{(\pi t)}]_0^{2.5}}{2.5\pi}=\frac{[a\sin{(\pi\times\frac{5}{2})}]}{2.5\pi}=\frac{a}{2.5\pi}
Adarsh
8 years ago
Sir what you mean by xavggt